OpenStudy (anonymous):

Can anyone help me with this question????????????????????????????????????????????????????? A math club had 7 members in its first year. It had 22 members in its fourth year and 32 members in its sixth year. Assuming a constant rate of increase, how many members will the math club have in its thirteenth year?

OpenStudy (ash2326):

First year club had 7 members Fourth year they have 22 Sixth year they have 32 so from first to fourth or in 3 years its members increase by (22-7)=15 from fourth to sixth or in 2 years its members increases by (32-22)=10 so in 3 years it increase by 15 \[or\ in\ a\ year\ members\ increase\ by\ 5 \] So in its thirteenth year no of years from first to thirteenth =(13-1)=12 years in 12 years members will increase by 12*5=60 In first year it has 7 members so in thirteenth year it will have=60+7=67 members

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

i was getting soooo confused on this

OpenStudy (ash2326):

Did you understand now?

OpenStudy (anonymous):

yesss i was getting confused because i didnt kn how to go from the fist all the way to the 13th